Presented here are the components in Wildhorse Co.s income statement. Determine the missing amounts. Sales Revenue Cost of GoOperating Expenses Net Income $ 20,840 (d) $13,360 $0 (e) $30,500 $47,810 $
Answer #1

--Requirement

Sales revenue Cost of Goods Sold Gross Profits Operating expenses Net Income
Year 1 $74,660 $40,460 $34,200 $20,840 $13,360
Year 2 $110,000 $74,400 $35,600 $5,100 $30,500
Year 3 $185,950 $75,920 $110,030 $47,810 $62,220

--Working calculation

Sales revenue Cost of Goods Sold Gross Profits Operating expenses Net Income
Year 1 74660 =74660-34200 34200 =34200-13360 13360
Year 2 110000 74400 =110000-74400 =35600-30500 30500
Year 3 =75920+110030 75920 110030 47810 =110030-47810

--Equations used:

Sales Revenue - Cost of Goods Sold = Gross Profits
Gross Profits - Operating expense = Net Income